What is the sound intensity level in the air near a jackhammer when the sound intensity level is 105 dB ?
Answer #1

Reference sound intensity Io = 10−12 W/m² (Threshold of hearing).

Sound intensity I = Io×10^(LI/10) W/m² = 10^−12×10^(105/10) W/m² = 0.0316 W/m².